OpenStudy (anonymous):

find the vertex of the function. write your answer if the form (x,y). Use the slash mark (/)as a fraction bar if necessary but do not use decimal points. y=x^2+10x+22

OpenStudy (anonymous):

help me ivettef365

OpenStudy (ivettef365):

x = -b/2a that is the formula

OpenStudy (cwrw238):

x^2+10x+22 = (x + 5)^2 - 25 + 22 = (x + 5)^2 - 3

OpenStudy (cwrw238):

yes thats formula for the x coordinate of the vertex = -10/2 = -5

OpenStudy (ivettef365):

right so x = -5, now find y

OpenStudy (cwrw238):

now you need the y-coordinate which you can get by plugging in -5 for x in the expression

OpenStudy (anonymous):

-3?

OpenStudy (cwrw238):

yep

OpenStudy (anonymous):

so -5,-3 answer ?

OpenStudy (ivettef365):

(-5)^2 +10(-5) + 22 = 25 -50 + 22 = 47 -50 = -3

OpenStudy (ivettef365):

correct

OpenStudy (cwrw238):

or you can get the vertex form converting to vertex form as i did above the x coordinate is value of x which makes parentheses = 0 ( that is -5) and the last part -3 = y coordinate